The width of a rectangle is 3 meters less than the length.The perimeter is 14 meters.Find the width.

The second angle of a triangle is twice the first angle.The third angle is 20
degrees more than the first angle.Find the first angle.

Answer by texttutoring(324)

Let w=width
Let l=length
The width is 3 metres less than the length:
w = l - 3
or, l = w+3

Perimeter of a rectangle is given by: P=2w+2l
Plug l=w+3 into the perimeter equation and solve for w:

P=2w+2l
P=2w + 2(w+3)
14 = 2w + 2w +6
8 = 4w
w = 2
The length is l = w+ 3 = 2+3 = 5.


Angle question:
Let a=first angle, b=2nd angle, c=3rd angle
b=2a
c=a+20

All triangles must have angle sums of 180 degrees, so:

a+b+c = 180

Now plug in b=2a and c=a+20 into this equation:

a+b+c=180
a + 2a + a+20 = 180
4a + 20 = 180
4a = 160
a = 40

The first angle is 40 degrees.